# CLoQ: Enhancing Fine-Tuning of Quantized LLMs via Calibrated LoRA Initialization

### v1

@@ Abstract
Fine-tuning large language models (LLMs) using low-rank adaptation (LoRA) has become a highly efficient approach for downstream tasks, particularly in scenarios with limited computational resources. However, applying LoRA techniques to quantized LLMs poses unique challenges due to the reduced representational precision of quantized weights. In this paper, we introduce CLoQ (Calibrated LoRA initialization for Quantized LLMs), a simplistic initialization strategy designed to overcome these challenges. Our approach focuses on minimizing the layer-wise discrepancy between the original LLM and its quantized counterpart with LoRA components during initialization. By leveraging a small calibration dataset, CLoQ quantizes a pre-trained LLM and determines the optimal LoRA components for each layer, ensuring a strong foundation for subsequent fine-tuning. A key contribution of this work is a novel theoretical result that enables the accurate and closed-form construction of these optimal LoRA components. We validate the efficacy of CLoQ across multiple tasks such as language generation, arithmetic reasoning, and commonsense reasoning, demonstrating that it consistently outperforms existing LoRA fine-tuning methods for quantized LLMs, especially at ultra low-bit widths. The code is available at https://github.com/AozhongZhang/CLoQ
@@ 1 Introduction
Large language models (LLMs) (Achiam et al., 2023; Touvron et al., 2023; Jiang et al., 2023; Guo et al., 2024a) have achieved remarkable success across a wide range of domains and applications. With ongoing advancements, the size and complexity of LLMs have grown exponentially, with some models now exceeding billions or even trillions of parameters. Although this scaling has unlocked unprecedented capabilities, it also introduces significant challenges, particularly in efficiently adapting these models to downstream tasks. Traditionally, full fine-tuning has been the dominant approach for adapting pre-trained models, involving updates to all model parameters. While effective in achieving state-of-the-art results, full fine-tuning is resource-intensive, requiring substantial GPU memory to store both model weights and optimizer states. These memory demands grow with the size of the model, making full fine-tuning increasingly impractical for large-scale models in resource-constrained settings.
@@
To address these challenges, parameter-efficient fine-tuning (PEFT) (Houlsby et al., 2019), such as Low-Rank Adaptation (LoRA) (Hu et al., 2021) , has emerged as a promising approach. PEFT updates only a small subset of parameters while keeping the majority of the model unchanged, enabling resource-efficient fine-tuning of large-scale models. LoRA, for instance, introduces small, learnable low-rank matrices into the model’s architecture. These matrices are optimized during fine-tuning while the original model weights remain frozen, significantly reducing memory and computational requirements. This design leverages the insight that weight updates often reside in a low dimensional subspace, allowing LoRA to achieve efficient adaptation with minimal overhead.
@@
In an orthogonal direction, model compression techniques , notably quantization (Rastegari et al., 2016; Hubara et al., 2018; Choi et al., 2018; Wang et al., 2018; Shao et al., 2023; Frantar et al., 2022a; Zhang et al., 2024a), have been developed to minimize GPU memory usage by converting high-precision weights into low-precision representations. Initially designed for inference in memory-limited environments, quantization methods have since been adapted to support fine-tuning. A notable advancement in this regard is QLoRA (Dettmers et al., 2023) , which combines LoRA with quantization techniques to reduce GPU memory requirements for fine-tuning significantly. By leveraging low-rank adaptation and quantized weights, QLoRA enables resource-efficient fine-tuning of LLMs without compromising performance, making it a powerful tool for adapting large-scale models to downstream tasks. However, extending LoRA to quantized LLMs introduces additional challenges, as the reduced representational precision of quantized weights can disrupt standard initialization strategies, impacting task performance. Recent works (Li et al., 2023; Yao et al., 2023; Liao et al., 2024; Guo et al., 2024b) proposed minimizing quantization error through strategic initialization of the low-rank components in LoRA, to align with the original weight states of the model. This strategy has demonstrated success in fine-tuning lower-bit quantized LLMs.
@@
Our CLoQ method requires no back-propagation, making it a highly efficient approach for fine-tuning quantized models. We demonstrate the effectiveness of CLoQ through extensive validation on multiple benchmark datasets. Our results show that CLoQ consistently outperforms existing LoRA methods for quantized LLMs, particularly at ultra-low bit-widths. For instance, the fine-tuning accuracy of INT2 CLoQ on the Llama2-13B model (Touvron et al., 2023) surpasses that of INT4 QLoRA (Dettmers et al., 2023) in the arithmetic reasoning tasks, as shown in Table 1.
@@ 2 Background
[…]
@@
Post-Training Quantization. Post-training quantization (PTQ) has been a cornerstone technique for compressing LLMs. PTQ methods directly identify low-precision representations of a model without requiring retraining, making them particularly well-suited for extremely large-scale AI models including LLMs. The simplest approach in this category is data-free quantization, commonly referred to as RTN, which involves rounding the pre-trained model weights to their nearest quantized states. More advanced PTQ algorithms, such as OPTQ (Frantar et al., 2022b), leverage a small calibration dataset to solve a least-squares problem under discrete constraints:
@@
to calibrate the quantization layer by layer. In this formulation, $\boldsymbol{X}$ denotes the activation or feature matrix corresponding to a batch of calibration data for a fixed layer. This approach ensures that the quantization process preserves the layer’s output by minimizing the discrepancy between the original and quantized representations on the calibration dataset. Several efficient back-propagation-free algorithms (Behdin et al., 2023; Zhang et al., 2024b) have been proposed to address (1).
Low-Rank Adaptation. LoRA (Hu et al., 2021) enables efficient fine-tuning of large pre-trained models by introducing two small, trainable matrices, $\boldsymbol{A}$ and $\boldsymbol{B}$, which are added to the frozen weight matrix $\boldsymbol{W}$. The weights of the fine-tuned model are then expressed as $\boldsymbol{W}+\boldsymbol{A}\boldsymbol{B}^{\top}$, where $\boldsymbol{A}\in\mathbb{R}^{m\times r}$, $\boldsymbol{B}\in\mathbb{R}^{n\times r}$, and $r\ll\min(m,n)$. During fine-tuning, only $\boldsymbol{A}$ and $\boldsymbol{B}$ are updated, while $\boldsymbol{W}$ remains fixed, significantly reducing the number of trainable parameters and computational overhead. LoRA initializes its parameters as follows:
@@
ensuring that the initialization maintains perfect alignment with the pre-trained weights $\boldsymbol{W}$. Recent research has focused on developing variants of LoRA aimed at enhancing its performance (Liu et al., 2024; Wang et al., 2024a, b).
@@ 3 Method
CLoQ is designed to enhance the fine-tuning of quantized LLMs by incorporating calibration data, building upon OPTQ (Frantar et al., 2022a) with a specific focus on the initialization of LoRA adapters. It utilizes second-order information derived from input activations $\boldsymbol{X}$ to ensure that the low-rank adapter matrices $\boldsymbol{A}$ and $\boldsymbol{B}$ are initialized in a manner that minimizes quantization error, particularly in relation to the data distribution. This alignment allows CLoQ to improve the fine-tuning process by making the low-rank adaptation more responsive to the model’s behavior during calibration, thereby reducing the mismatch between the quantized model and its full-precision counterpart. Importantly, CLoQ uses the same calibration data as OPTQ, ensuring that the initialization of LoRA adapters is both universal and effective across various downstream tasks. In our experiments, we use the Wikitext dataset for calibration, offering a robust and generalizable foundation for fine-tuning across a range of task domains.
@@ 3.1 Calibrated Quantization and Low-Rank Initialization
This problem can, in theory, be solved using an alternating minimization method, whose $t$-th iteration reads:
@@ 3.1.1 Post-Training Quantization
an optimization problem that has been extensively studied in recent literature (Frantar et al., 2022b; Chee et al., 2023; Xiao et al., 2023; Zhang et al., 2024b). For this, we adopt the widely-used OPTQ method (Frantar et al., 2022a). To further enhance OPTQ’s performance, we incorporate a preprocessing technique called weight magnitude reduction (MagR) (Zhang et al., 2024a), which modifies $\boldsymbol{W}$ by removing outliers prior to quantization. MagR preprocessing significantly improves OPTQ’s effectiveness in the low-bit regime while introducing minimal additional time beyond the OPTQ process and incurring no computational or memory overhead during inference time.
@@ 3.1.2 Generalized Low-rank Approximation
Based on these facts, we interpret problem (4) as finding the standard best rank-$r$ approximation of $\boldsymbol{R}\Delta\boldsymbol{W}$, $\mathrm{LR}_{r}(\boldsymbol{R}\Delta\boldsymbol{W})$, which can be obtained by performing the SVD of $\boldsymbol{R}\Delta\boldsymbol{W}$ and extracting the top-$r$ principal components. Then, $(\boldsymbol{A},\boldsymbol{B})$ is an optimal solution to
@@
  • One SVD is required to compute $\boldsymbol{R}$ and another is needed to determine $\mathrm{LR}_{r}(\boldsymbol{R}\Delta\boldsymbol{W})$. Given that $\boldsymbol{R}\in\mathbb{R}^{m\times m}$ and $\boldsymbol{R}\Delta\boldsymbol{W}\in\mathbb{R}^{m\times n}$, while $\boldsymbol{X}\in\mathbb{R}^{(b\cdot l)\times m}$, the computational complexity of SVDs required for solving (4) is independent of $b\cdot l$, which is significantly larger than $m$ or $n$ in practice. Here $l$ represents the context length, and $b$ denotes the size of the calibration dataset.
  • Indeed, (5) admits infinitely many optimal solutions. Suppose $\mathrm{LR}_{r}(\boldsymbol{R}\Delta\boldsymbol{W})$ has the form $\boldsymbol{U}_{:r}\boldsymbol{\Sigma}_{:r}\boldsymbol{V}_{:r}^{\top}$. In our experiments, we consistently take $\boldsymbol{A}=\boldsymbol{R}^{-1}\boldsymbol{U}_{:r}\boldsymbol{\Sigma}_{:r}$ and $\boldsymbol{B}=\boldsymbol{V}_{:r}$, which empirically performs well. However, it is clear that if $(\boldsymbol{A},\boldsymbol{B})$ is an optimal solution, then for any invertible $\boldsymbol{C}\in\mathbb{R}^{r\times r}$, the pair $\left(\boldsymbol{A}\boldsymbol{C},\,\boldsymbol{B}(\boldsymbol{C}^{-1})^{\top}\right)$ also satisfies (5) and thus provides an optimal solution to (4). For example, $(\boldsymbol{A},\boldsymbol{B})=(\boldsymbol{R}^{-1}\boldsymbol{U}_{:r},\,\boldsymbol{V}_{:r}\boldsymbol{\Sigma}_{:r})$ or $(\boldsymbol{R}^{-1}\boldsymbol{U}_{:r}\boldsymbol{\Sigma}_{:r}^{\frac{1}{2}},\,\boldsymbol{V}_{:r}\boldsymbol{\Sigma}_{:r}^{\frac{1}{2}})$. In Section 5, our ablation study shows that the combination $(\boldsymbol{R}^{-1}\boldsymbol{U}_{:r}\boldsymbol{\Sigma}_{:r},\boldsymbol{V}_{:r})$ gives the best practical performance during the subsequent LoRA fine-tuning process. A more comprehensive theoretical analysis of different initialization schemes (Hayou et al., 2024) will be addressed in future work.
  • When $\boldsymbol{H}$ is not invertible or poorly conditioned, we propose adding a small constant $\lambda$ to the diagonal elements. Specifically, we typically set $\lambda=0.01\frac{\mathrm{Tr}(\boldsymbol{H})}{m}$, following a strategy similar to that used in the prior works (Frantar et al., 2022b; Chee et al., 2023). This adjustment has consistently proven effective in mitigating numerical issues and ensuring stability during computations.
@@
Input: Gram matrix of activations $\boldsymbol{H}=\boldsymbol{X}^{\top}\boldsymbol{X}\in\mathbb{R}^{m\times m}$, Pre-trained weight matrix $\boldsymbol{W}\in\mathbb{R}^{m\times n}$, Rank $r\ll\min(m,n)$.

### v2

@@ Abstract
Fine-tuning large language models (LLMs) using low-rank adaptation (LoRA) has become a highly efficient approach for downstream tasks, particularly in scenarios with limited computational resources. However, applying LoRA techniques to quantized LLMs poses unique challenges due to the reduced representational precision of quantized weights. In this paper, we introduce CLoQ (Calibrated LoRA initialization for Quantized LLMs), a simplistic initialization strategy designed to overcome these challenges. Our approach focuses on minimizing the layer-wise discrepancy between the original LLM and its quantized counterpart with LoRA components during initialization. By leveraging a small calibration dataset, CLoQ quantizes a pre-trained LLM and determines the optimal LoRA components for each layer, ensuring a strong foundation for subsequent fine-tuning. A key contribution of this work is a novel theoretical result that enables the accurate and closed-form construction of these optimal LoRA components. We validate the efficacy of CLoQ across multiple tasks such as language generation, arithmetic reasoning, and commonsense reasoning, demonstrating that it consistently outperforms existing LoRA fine-tuning methods for quantized LLMs, especially at 2-bit. The code is available at https://github.com/AozhongZhang/CLoQ
@@ 1 Introduction
Large language models (LLMs) Achiam et al. (2023); Touvron et al. (2023); Jiang et al. (2023); Guo et al. (2024a) have achieved remarkable success across a wide range of domains and applications. With ongoing advancements, the size and complexity of LLMs have grown exponentially, with some models now exceeding billions or even trillions of parameters. Although this scaling has unlocked unprecedented capabilities, it also introduces significant challenges, particularly in efficiently adapting these models to downstream tasks. Traditionally, full fine-tuning has been the dominant approach for adapting pre-trained models, involving updates to all model parameters. While effective in achieving state-of-the-art results, full fine-tuning is resource-intensive, requiring substantial GPU memory to store both model weights and optimizer states. These memory demands grow with the size of the model, making full fine-tuning increasingly impractical for large-scale models in resource-constrained settings.
@@
To address these challenges, parameter-efficient fine-tuning (PEFT) Houlsby et al. (2019), such as Low-Rank Adaptation (LoRA) Hu et al. (2021) , has emerged as a promising approach. PEFT updates only a small subset of parameters while keeping the majority of the model unchanged, enabling resource-efficient fine-tuning of large-scale models. LoRA, for instance, introduces small, learnable low-rank matrices into the model’s architecture. These matrices are optimized during fine-tuning while the original model weights remain frozen, significantly reducing memory and computational requirements. This design leverages the insight that weight updates often reside in a low dimensional subspace, allowing LoRA to achieve efficient adaptation with minimal overhead.
@@
In an orthogonal direction, model compression techniques , notably quantization Rastegari et al. (2016); Hubara et al. (2018); Choi et al. (2018); Wang et al. (2018); Yin et al. (2016; 2018; 2019b; 2019a); Li et al. (2023b); Shao et al. (2023); Zhang et al. (2023); Frantar et al. (2022a); Zhang et al. (2024a), have been developed to minimize GPU memory usage by converting high-precision weights into low-precision representations. Initially designed for inference in memory-limited environments, quantization methods have since been adapted to support fine-tuning. A notable advancement in this regard is QLoRA Dettmers et al. (2023) , which combines LoRA with quantization techniques to reduce GPU memory requirements for fine-tuning significantly. By leveraging low-rank adaptation and quantized weights, QLoRA enables resource-efficient fine-tuning of LLMs without compromising performance, making it a powerful tool for adapting large-scale models to downstream tasks. However, extending LoRA to quantized LLMs introduces additional challenges, as the reduced representational precision of quantized weights can disrupt standard initialization strategies, impacting task performance. Recent works Li et al. (2023a); Yao et al. (2023); Liao et al. (2024); Guo et al. (2024b) proposed minimizing quantization error through strategic initialization of the low-rank components in LoRA, to align with the original weight states of the model. This strategy has demonstrated success in fine-tuning lower-bit quantized LLMs.
@@
Our CLoQ method requires no back-propagation, making it a highly efficient LoRA initialization scheme for fine-tuning quantized models. We demonstrate the effectiveness of CLoQ through extensive validation on multiple benchmark datasets. Our results show that CLoQ consistently outperforms existing LoRA methods for quantized LLMs, particularly at ultra-low bit-widths. For instance, the fine-tuning accuracy of INT2 CLoQ on the Llama2-13B model Touvron et al. (2023) surpasses that of INT4 QLoRA Dettmers et al. (2023) in the arithmetic reasoning tasks, as shown in Table 3.
@@ 2 Background
Integer Quantizer. Given a set of $m$ weights $\boldsymbol{w}\in\mathbb{R}^{m}$, the widely-used $b$-bit uniform integer (INT) quantizer Choi et al. (2018) determines the float scaling factor $\delta=\frac{\max(\boldsymbol{w})-\min(\boldsymbol{w})}{2^{b}-1}$ and zero-point $z=-\left\lfloor\frac{\min(\boldsymbol{w})}{\delta}\right\rceil$, where $\left\lfloor\cdot\right\rceil$ is nearest-to-round operation. The quantizer projects $\boldsymbol{w}$ onto the equally spaced grids $\mathbb{Q}=\{z\cdot\delta,(z+1)\cdot\delta,\ldots,\left(z+(2^{b}-1)\right)\cdot\delta\}^{m}$ to obtain the quantized weights
@@
Post-Training Quantization. Post-training quantization (PTQ) has been a cornerstone technique for compressing LLMs. PTQ methods directly identify low-precision representations of a model without requiring retraining, making them particularly well-suited for extremely large-scale AI models including LLMs. The simplest approach in this category is data-free quantization, commonly referred to as RTN, which involves rounding the pre-trained model weights to their nearest quantized states. More advanced PTQ algorithms, such as OPTQ Frantar et al. (2022b), leverage a small calibration dataset to solve a least-squares problem under discrete constraints:
@@
to calibrate the quantization layer by layer, where $\boldsymbol{W}$ is the weight matrix of the pre-trained model. In this formulation, $\boldsymbol{X}$ denotes the activation or feature matrix corresponding to a batch of calibration data for a fixed layer. This approach ensures that the quantization process preserves the layer’s output by minimizing the discrepancy between the original and quantized representations on the calibration dataset. Several efficient back-propagation-free algorithms Zhang et al. (2023); Behdin et al. (2023); Zhang et al. (2024b) have been proposed to address (1).
Low-Rank Adaptation. LoRA Hu et al. (2021) enables efficient fine-tuning of large pre-trained models by introducing two small, trainable matrices, $\boldsymbol{A}$ and $\boldsymbol{B}$, which are added to the frozen weight matrix $\boldsymbol{W}$. The weights of the fine-tuned model are then expressed as $\boldsymbol{W}+\boldsymbol{A}\boldsymbol{B}^{\top}$, where $\boldsymbol{A}\in\mathbb{R}^{m\times r}$, $\boldsymbol{B}\in\mathbb{R}^{n\times r}$, and $r\ll\min(m,n)$. During fine-tuning, only $\boldsymbol{A}$ and $\boldsymbol{B}$ are updated, while $\boldsymbol{W}$ remains fixed, significantly reducing the number of trainable parameters and computational overhead. LoRA initializes its parameters as follows:
@@
ensuring that the initialization maintains perfect alignment with the pre-trained weights $\boldsymbol{W}$. Recent research has focused on developing variants of LoRA aimed at enhancing its performance Wang et al. (2024a; b); Liu et al. (2024); Wang et al. (2024c); Meng et al. (2024); Büyükakyüz (2024).
@@ 3 Method
CLoQ is designed to enhance the fine-tuning of quantized LLMs by incorporating calibration data, building upon OPTQ Frantar et al. (2022a) with a specific focus on the initialization of LoRA adapters. It utilizes second-order information derived from input activations $\boldsymbol{X}$ to ensure that the low-rank adapter matrices $\boldsymbol{A}$ and $\boldsymbol{B}$ are initialized in a manner that minimizes quantization error, particularly in relation to the data distribution. This alignment allows CLoQ to improve the fine-tuning process by making the low-rank adaptation more responsive to the model’s behavior during calibration, thereby reducing the mismatch between the quantized model and its full-precision counterpart. Importantly, CLoQ uses the same calibration data as OPTQ, ensuring that the initialization of LoRA adapters is both universal and effective across various downstream tasks. In our experiments, we use the Wikitext dataset for calibration, offering a robust and generalizable foundation for fine-tuning across a range of task domains.
@@ 3.1 Calibrated Quantization and Low-Rank Initialization
This problem can, in theory, be solved using an alternating minimization (AltMin) method, whose $t$-th iteration reads:
@@ 3.1.1 Post-Training Quantization
an optimization problem that has been extensively studied in recent literature Frantar et al. (2022b); Zhang et al. (2023); Chee et al. (2023); Xiao et al. (2023); Zhang et al. (2024b). For this, we adopt the widely-used OPTQ method Frantar et al. (2022a). To further enhance OPTQ’s performance, we incorporate a preprocessing technique called weight magnitude reduction (MagR) Zhang et al. (2024a), which modifies $\boldsymbol{W}$ by removing outliers prior to quantization. MagR preprocessing significantly improves OPTQ’s effectiveness in the low-bit regime while introducing minimal additional time beyond the OPTQ process and incurring no computational or memory overhead during inference time.
@@ 3.1.2 Generalized Low-rank Approximation
Based on these facts, we interpret problem (4) as finding the standard best rank-$r$ approximation of $\boldsymbol{R}\Delta\boldsymbol{W}$, $\mathrm{LR}_{r}(\boldsymbol{R}\Delta\boldsymbol{W})$, which can be obtained by performing the SVD of $\boldsymbol{R}\Delta\boldsymbol{W}$ and extracting the top-$r$ principal components Eckart & Young (1936). Then, $(\boldsymbol{A},\boldsymbol{B})$ is an optimal solution to
@@
  • One SVD is required to compute $\boldsymbol{R}$ and another is needed to determine $\mathrm{LR}_{r}(\boldsymbol{R}\Delta\boldsymbol{W})$. Given that $\boldsymbol{R}\in\mathbb{R}^{m\times m}$ and $\boldsymbol{R}\Delta\boldsymbol{W}\in\mathbb{R}^{m\times n}$, while $\boldsymbol{X}\in\mathbb{R}^{(b\cdot l)\times m}$, the computational complexity of SVDs required for solving (4) is independent of $b\cdot l$, which is significantly larger than $m$ or $n$ in practice. Here $l$ represents the context length, and $b$ denotes the size of the calibration dataset. We note that CLoQ requires fewer SVD computations than LoftQ Li et al. (2023a), which, by default, performs five AltMin iterations, each involving one SVD. Although CLoQ uses the more expensive GPTQ for the quantization (versus the data-free RTN in LoftQ), their overall initialization runtimes are comparable, as shown in Table 10.
  • Indeed, (5) admits infinitely many optimal solutions. Suppose $\mathrm{LR}_{r}(\boldsymbol{R}\Delta\boldsymbol{W})$ has the form $\boldsymbol{U}_{:r}\boldsymbol{\Sigma}_{:r}\boldsymbol{V}_{:r}^{\top}$. In our experiments, we consistently take $\boldsymbol{A}=\boldsymbol{R}^{-1}\boldsymbol{U}_{:r}\boldsymbol{\Sigma}_{:r}$ and $\boldsymbol{B}=\boldsymbol{V}_{:r}$, which empirically performs well. However, it is clear that if $(\boldsymbol{A},\boldsymbol{B})$ is an optimal solution, then for any invertible $\boldsymbol{C}\in\mathbb{R}^{r\times r}$, the pair $\left(\boldsymbol{A}\boldsymbol{C},\,\boldsymbol{B}(\boldsymbol{C}^{-1})^{\top}\right)$ also satisfies (5) and thus provides an optimal solution to (4). For example, $(\boldsymbol{A},\boldsymbol{B})=(\boldsymbol{R}^{-1}\boldsymbol{U}_{:r},\,\boldsymbol{V}_{:r}\boldsymbol{\Sigma}_{:r})$ or $(\boldsymbol{R}^{-1}\boldsymbol{U}_{:r}\boldsymbol{\Sigma}_{:r}^{\frac{1}{2}},\,\boldsymbol{V}_{:r}\boldsymbol{\Sigma}_{:r}^{\frac{1}{2}})$. In Section 5, our ablation study shows that the combination $(\boldsymbol{R}^{-1}\boldsymbol{U}_{:r}\boldsymbol{\Sigma}_{:r},\boldsymbol{V}_{:r})$ gives the best practical performance during the subsequent LoRA fine-tuning process. A more comprehensive theoretical analysis of the impact of initialization like Hayou et al. (2024); Li et al. (2024) will be addressed in future work.
  • When $\boldsymbol{H}$ is not invertible or poorly conditioned, we propose adding a small constant $\lambda$ to the diagonal elements. Specifically, we typically set $\lambda=0.01\frac{\mathrm{Tr}(\boldsymbol{H})}{m}$, following a strategy similar to that used in the prior works Frantar et al. (2022b); Chee et al. (2023). This adjustment has consistently proven effective in mitigating numerical issues and ensuring stability during computations.
  • In theory, even if $\boldsymbol{X}$ is rank-deficient, the optimality condition $\boldsymbol{R}\boldsymbol{A}\boldsymbol{B}^{\top}=\mathrm{LR}_{r}(\boldsymbol{R}\Delta\boldsymbol{W})$ in the proof of Theorem 3.1 still holds. But in this case, since $\boldsymbol{R}$ is also rank-deficient, this optimality condition permits infinitely many solution for $\boldsymbol{A}\boldsymbol{B}^{\top}$, and we may simply take $\boldsymbol{A}\boldsymbol{B}^{\top}=\boldsymbol{R}^{\dagger}\mathrm{LR}_{r}(\boldsymbol{R}\Delta\boldsymbol{W})$, where $\boldsymbol{R}^{\dagger}$ is a pseudo-inverse.
@@
Input: Regularized Gram matrix of activations $\boldsymbol{H}=\boldsymbol{X}^{\top}\boldsymbol{X}+\lambda\boldsymbol{I}\in\mathbb{R}^{m\times m}$, Pre-trained weight matrix $\boldsymbol{W}\in\mathbb{R}^{m\times n}$, Rank $r\ll\min(m,n)$.
@@
4 Experiment
Models and Datasets. We test CLoQ on Llama2-7b, Llama2-13b Touvron et al. (2023), Llama3-8b Grattafiori et al. (2024) and Mistral-7b-v0.1 Jiang et al. (2023) models. Following prior works Frantar et al. (2022a), we randomly sample 128 instances, each with a context length of 2048 tokens, from the WikiText-2 dataset Merity et al. (2016) to serve as the calibration set for quantization. Then, we fine-tune and evaluate the models on WikiText-2 for language modeling. For single arithmetic reasoning tasks, we fine-tune and evaluate on the GSM8K Cobbe et al. (2021). For multi arithmetic reasoning, we fine-tune the models on Math10K Hu et al. (2023) and then evaluate the test sets of AQuA Ling et al. (2017), GSM8K, MAWPS Koncel-Kedziorski et al. (2016) and SVAMP Patel et al. (2021). For commonsense reasoning tasks, we fine-tune the models on Commonsense170K Hu et al. (2023) and evaluate on eight representative tasks: BoolQ Clark et al. (2019), PIQA Bisk et al. (2020), SIQA Sap et al. (2019), HellaSwag Zellers et al. (2019), WinoGrande Sakaguchi et al. (2021), ARC-e, ARC-c Clark et al. (2018) and OBQA Mihaylov et al. (2018).
Baselines. We compare with LoRA Hu et al. (2021), QLoRA Dettmers et al. (2023), GPTQ-LoRA GPT (2023), LoftQ Li et al. (2023a) and ApiQ Liao et al. (2024). LoRA is often considered as the benchmark for fine-tuning performance. QLoRA incorporates NF-quantization, with its low-rank initialization aligning with the standard LoRA method. GPTQ-LoRA integrates OPTQ for the base weights and fine-tunes the LoRA component while preserving its low-rank initialization as in the standard LoRA, with the quantized weights kept frozen. In contrast, LoftQ and LQ-LoRA carefully initialize the quantized weight and low-rank matrices by solving some optimization problems to minimize the approximation error. Furthermore, ApiQ uses gradient-based block-wise optimization to specifically initialize the low-rank components during post-training quantization.
4.1 Implementation details
Quantization. We quantize the weights of all linear layers in the base model using MagR preprocessing Zhang et al. (2024a) followed by OPTQ Frantar et al. (2022a). The quantization scheme employs uniform (a.k.a. INT) and asymmetric quantization, with a default group size of 64. After quantization, we compute the LoRA components $\boldsymbol{A}$ and $\boldsymbol{B}$, which are maintained in FP16 precision.
Fine-tuning. Following prior works Dettmers et al. (2023); Li et al. (2023a); Liao et al. (2024), we fine-tune the models using the standard LoRA configuration, with modifications to the LoRA initialization and learning rates. The quantized weights remain fixed, and only the LoRA adapter matrices are trainable during fine-tuning. The rank of the LoRA adapters is consistently set to 64 across all methods. For optimization, we use AdamW Loshchilov (2017). All experiments are conducted on NVIDIA A100 GPUs with 80GB of memory.
4.2 Fine-tuning Results
Language modeling. We evaluate the models by reporting the perplexity of language generation on WikiText-2. As shown in Table 1 and 2, CLoQ consistently outperforms existing methods across most bit levels and model architectures. Notably, at INT2, CLoQ proves effective, achieving a perplexity reduction of 0.95 over ApiQ-lw and 1.34 over LoftQ on Llama2-7B, and even surpasses ApiQ-bw by 0.1 perplexity. This result highlights CLoQ’s ability to maintain superior performance even under ultra-low bit quantization constraints.
Llama2-7B Llama2-13B
[…]
LoRA 16 5.08 36.9 5.12 45.3
QLoRA 4 5.70 35.1 5.22 39.9
LoftQ 4 5.24 35.0 5.16 45.0
ApiQ-lw 4 5.28 36.4 4.78 50.4
ApiQ-bw 4 5.27 39.8 4.78 51.2
CLoQ 4 5.25 40.6 4.78 49.3
QLoRA 3 5.73 32.1 5.22 40.7
LoftQ 3 5.63 32.9 5.13 44.4
ApiQ-lw 3 5.53 36.0 4.98 45.4
ApiQ-bw 3 5.49 39.3 4.96 47.6
CLoQ 3 5.45 39.9 4.92 48.1
QLoRA 2 N.A. N.A. N.A. N.A.
LoftQ 2 7.85 20.9 7.69 25.4
ApiQ-lw 2 7.46 26.0 6.29 36.3
ApiQ-bw 2 6.61 33.5 5.79 41.2
CLoQ 2 6.51 33.7 5.73 41.7
Table 2: Finetuning results of WikiText and GSM8K on Llama3-8B and Mistral-7B.
Llama3-8B Mistral-7B
[…]
LoRA 16 6.34 47.8 5.17 52.8
LoftQ 2 14.09 31.6 1849.33 1.7
ApiQ-lw 2 - - 7.18 41.3
ApiQ-bw 2 9.89 44.4 6.69 45.0
CLoQ 2 9.49 45.4 6.68 45.0
Arithmetic reasoning (single task). To evaluate the models on GSM8K, we extract numerical answers from the generated solutions and determine accuracy by analyzing these extracted values. As shown in Table 1 and 2, CLoQ achieves better performance across different model types and quantization bit levels. At INT2, CLoQ reaches an accuracy of $33.7\%$ on Llama2-7B, which CLoQ achieves a $7.7\%$ improvement in performance compared with ApiQ-lw. Moreover, on the Llama3-8B model, CLoQ even outperforms ApiQ-bw by $1.0\%$ at INT2. More remarkably, CLoQ achieves comprehensive superiority over all existing methods at INT2 and INT3 on Llama2-13B. On the Mistral-7B model, the accuracy of INT2 CLoQ is better than ApiQ-lw and is comparable with ApiQ-bw.
Arithmetic reasoning. To evaluate CLoQ’s effectiveness across multiple arithmetic reasoning tasks, we fine-tuned models on the Math10K dataset and assessed their performance on four separate math reasoning benchmarks, demonstrating their adaptability to diverse mathematical challenges. As shown in Table 3 and 4, CLoQ consistently outperforms other methods across various model architectures and quantization levels, achieving superior accuracy both on average and on individual datasets. Notably, at INT2 quantization, CLoQ delivers substantial gains, surpassing ApiQ-bw by $1.9\%$ on LLaMA2-7B and by $2.1\%$ on both LLaMA3-8B and LLaMA2-13B. Even at INT4, CLoQ surpasses both LoRA and QLoRA on Llama2-7B. Moreover, for complex reasoning tasks such as GSM8K and SVAMP, CLoQ achieves 4.1% and 1.7% higher accuracy than ApiQ-bw at INT2 on Llama3-8B, highlighting its robustness in challenging problem settings. These results underscore CLoQ’s remarkable potential in handling intricate reasoning tasks with enhanced accuracy, even under ultra-low-bit quantization.